OpenStudy (anonymous):

Find the general solution: 9y''-12y'+4y=0

OpenStudy (anonymous):

\[9m^2-12m+4=0\]

OpenStudy (anonymous):

can u factorize this polynomial ?

OpenStudy (anonymous):

(3m+2)(3m+2)

OpenStudy (anonymous):

@mustafa2014

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(3m-2)(3m-2)=0

OpenStudy (anonymous):

minus not plus

OpenStudy (anonymous):

3m=2 >>> m=2/3 , m=2/3

OpenStudy (anonymous):

\[y _{c}=C _{1}+xC _{2}e ^{mx}\]

OpenStudy (anonymous):

y = C1+xC2e^(2/3)x ?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[y _{c}=C _{1}e ^{mx}+xC _{2}e ^{mx}\]

OpenStudy (anonymous):

C1 times e^mx

OpenStudy (anonymous):

y=C1e^(2/3)x + xC2e^(2/3)x

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

Thank you, can we do another?